OpenStudy (anonymous):

Choose the correct simplification of the expression (5x^4y^2z^2)(3x^4y^3z^5). 15x^16y^6z^10 15x^8y^5z^7 8x^8y^5z^7 8x^16y^6z^10

OpenStudy (mathstudent55):

Multiply the numbers together: 3 * 5 = 15 Multiply similar variables together using x^mx^n = x^(m + n) For example, x^3 * x*5 = x^(3 + 5) = x^8 If you see a variable without an exponent, then use 1 as its exponent. For example, x * x^3 = x^1 * x^3 = x^4 x^4 * x^4 = x^8 (add the exponents) y^2 * y^3 = y^5 z^2 * z^5 = z^7